OpenStudy (anonymous):

If the sphere shown above has a radius of 5 units, then what is the approximate volume of the sphere? 261.67 cubic units 314 cubic units 523.33 cubic units

OpenStudy (anonymous):



OpenStudy (jdoe0001):

\(\Large \textit{volume of a sphere}=\cfrac{4}{3}\pi r^3\qquad r=radius\)

OpenStudy (anonymous):

sorry man am still confused am i supposed to plug in 5 @jdoe0001

OpenStudy (jdoe0001):

" a radius of 5 units," \( \textit{volume of a sphere}=\cfrac{4}{3}\pi r^3\qquad r=radius=5 \implies \cfrac{4}{3}\pi 5^3\implies ?\)